# GOFA: A Generative One-For-All Model for Joint Graph Language Modeling

### v1

@@ Appendix D Related Work Extended
GNNs and Transformers: In recent years, GNNs have become the most popular method for dealing with graph learning problems due to their extraordinary ability in structural learning. Particularly, Previous works Xu et al. [2018], Morris et al. [2019] show that the expressive power of message-passing GNNs can be as powerful as the 1-dimensional Weisfeiler-Lehman test, a powerful algorithm for graph isomorphism problems. Many recent works also try to design more powerful GNNs that beyond the 1-WL test Zhang and Li [2021], Kong et al. [2022], Feng et al. [2022], Huang et al. [2023b], Zhang et al. [2023b, a], Feng et al. [2023], Kong et al. [2023] for better structural ability like learning distance between nodes or counting cycles in graph. Some works try to combine the GNN with the transformer. particularly, GraphFormers Yang et al. [2021] and GROVER Rong et al. [2020] also insert a GNN layer between consecutive transformer layers for modeling graph inductive bias. Different from us, their transformer layers are randomly initialized and directly tuned on downstream tasks without text.
Appendix E Graph Structure Question Example of LLM
@@ F.1 General settings
Subgraph sampling: In the GOFA, for node/link/graph-level tasks, the input format is unified as a subgraph task. Namely, for node/link-level tasks, we will select a $k$-hop subgraph surrounding the target nodes as the input graph for the model. We follow a similar subgraph sampling strategy as OFA Liu et al. [2023a] with minor improvement. Specifically, for node-level tasks, we directly sample the $k$-hop subgraph rooted at the target node. Meanwhile, we set a threshold for maximum nodes per hop. If the nodes in a certain hop exceed the threshold, we will randomly sample maximum nodes from all nodes. For link-level tasks, we will independently do the sampling for both two nodes and then merge two subgraphs as the final subgraph.
@@
Implementations. Both the GOFA and all baselines are implemented using Python with Pytorch, transformers, and PyG Fey and Lenssen [2019] packages.
F.2 Pre-training of GOFA
@@
Dataset. For pretraining, we use two datasets: MAG240M and Ultrachat200k. For MAG240M, each training sample is a subgraph sampled around a node. Next, we will construct sentence completion, Shortest path distance, and common neighbor tasks for each sample. Specifically, given a sampled subgraph rooted at the target node, we will first build the sentence completion task on the target node (the root node in the sampled subgraph). Next, we randomly sample $k$ nodes in the subgraph. Next, for each node, we will independently construct all three tasks. For the shortest path distance and common neighbor, the task is to return the results between the selected nodes and the target node. In the pretraining, we set $k$ as 3, which results in 10 tasks for each training sample. For Ultrachat200k, we directly use the constructed graphs for pertaining. The detailed task prompts and answer examples are shown in Table 7.
@@
Training details. The initial weight of the LLM compressor and decoder is obtained from ICAE Ge et al. [2023]. The initial weight of all GNN layers is randomly initialized. The value of all gates in the residual connection is set to 0 to ensure the initialized model performs the same as the original language model. During the training, we only tune the GNN layers. Due to limited computation resources, we only include 1500000 training samples from MAG240M and 120000 training samples from Ultrachat200k. We pre-train two models using Llama2-7b and Mistral-7b, respectively. The training is conducted on 8 NVIDIAA100_SXM4_80GB GPUs with DeepSpeed stage 2 Rajbhandari et al. [2020] parallelism. The detailed training parameters are set the same for both two models and are listed in Table 8. We use AdamW optimizer with $\beta=(0.9,0.95)$. We use a cosine annealing learning rate scheduler, and the minimum learning rate is 10% of the initial learning rate. We restarted the learning rate 2 times on one-third and two-thirds of the training.
@@
Table 9: Prompt examples of GOFA for each dataset in Zero-shot learning.
@@
Cora-node You need to answer this question based on the information from this node: [NODEID]. This is a co-citation network focusing on artificial intelligence, nodes represent academic papers and edges represent two papers that are co-cited by other papers. You are an expert in computer science. You need to choose the correct paper category based on the paper content and its co-citation network. For example, if the paper [NODEID] {<label_description>, choose <label>;}. What is the most likely paper category for the target paper? Choose from the following: {<label>}.
@@
Cora-link You need to answer this question based on the information from this node: [NODEID1] and [NODEID2]. This is a co-citation network focusing on artificial intelligence, nodes represent academic papers, and edges represent two papers that are co-cited by other papers. You are a computer science expert tasked with determining whether two given papers are co-cited by another paper based on their content and network characteristics. Evaluate the following criteria: assess whether the topics of the two papers are similar, check if the shortest path distance between the two papers is small, and verify whether the papers have a large number of common neighbors in the citation network. If the answer to most of these questions is Yes, choose Yes; if the answer to most of these questions is No, choose No.
@@
PubMed-node You need to answer this question based on the information from this node: [NODEID]. This is a co-citation network from the Pubmed platform focusing on diabetes mellitus. Nodes represent academic papers and edges represent two papers that are co-cited by other papers. You are an expert on diabetes mellitus. You need to choose the correct paper category based on the paper content and its co-citation network. For example, if the paper [NODEID] {<label_description>, choose <label>;}. What is the most likely paper category for the target paper? Choose from the following: {<label>}.
WikiCS You need to answer this question based on the information from this node: [NODEID]. This is a Wikipedia graph focusing on computer science. Nodes represent Wikipedia terms and edges represent two terms that have hyperlinks. You are an expert in computer science. You need to choose the correct category of Wikipedia term based on the term content. For example, if the term [NODEID] {<label_description>, choose <label>;}. What is the most like category for this Wikipedia entry? Choose from the following: {<label>}.
Products You need to answer this question based on the information from this node: [NODEID]. This is a co-purchase network from the Amazon platform. Nodes represent the products sold on Amazon and edges represent two products that are co-purchased together. For example, if the product [NODEID] {<label_description>, choose <label>;}. What is the most like category for this product? Choose from the following: {<label>}.
@@
FB15K237 You need to answer this question based on the information from this node: [NODEID1] and [NODEID2]. This is a knowledge graph from the FreeBase. Nodes represent knowledge entities and edges represent relations between two entities. You are an expert in knowledge graph reasoning. You need to choose the correct relation type between two target entities based on their existing relations. For example, if two relations {<label_description>, choose <label>;}. What is the relationship between two target entities? Choose from the following list: {<label>}."
@@
F.3 Zero-shot learning
Setting. For the zero-shot learning, we select Cora-link, Cora-node, PubMed-node, WikiCS, Products, ExplaGraphs, and SceneGraphs as evaluation datasets. For all datasets, we directly evaluate baselines and GOFA on the test set.
Detail of Baselines. For baseline methods, we compare the GOFA with two types of methods. The first type of method is to evaluate the performance directly using the LLM models. We select Llama2-7B and Mistral-7B Jiang et al. [2023] as baseline models. For these models, we directly give the well-trained model the content in all target nodes and then concatenate the same prompt we use in the GOFA. The second type of method is the graph foundation model. We select the OFA Liu et al. [2023a] and GraphGPT Tang et al. [2023] as baseline methods. For OFA, we directly run the code from the original source to train the model using the Arxiv and FB15K237 datasets. We train the model 50 epochs and all other settings following the default OFA setting. To ensure a fair comparison, we use the Llama2-7b as the embedding model for OFA and report the test performance. For GraphGPT, we directly report the result from the original paper.

### v2

@@ Appendix D Related Work Extended
Prompt Learning in Graph: The success of foundation models inspired many works to adapt their power to the graph domain. Earlier attempts designed a graph prompting mechanism such that a trained model can adapt to new data by fine-tuning a soft prompting vector (Liu et al., 2023c; Yu et al., 2023; Sun et al., 2023; Xia et al., 2024). GraphPrompt (Liu et al., 2023c; Yu et al., 2023) pretrains on link prediction tasks, and then finetune a prompt matrix for each downstream node or graph classification task. All in One (Sun et al., 2023) designs prompt tokens that are used to modify node features and then take a meta-learning paradigm for multi-task learning. Subsequent works extend graph prompts to allow in-context learning without weight update (Huang et al., 2023a; Galkin et al., 2023). However, these works only tackle limited types of tasks and do not generalize to new domains. Hence, researchers propose integrating LLM into the graph learning.
GNNs and Transformers: In recent years, GNNs have become the most popular method for dealing with graph learning problems due to their extraordinary ability in structural learning. Particularly, Previous works (Xu et al., 2018; Morris et al., 2019) show that the expressive power of message-passing GNNs can be as powerful as the 1-dimensional Weisfeiler-Lehman test, a powerful algorithm for graph isomorphism problems. Many recent works also try to design more powerful GNNs that beyond the 1-WL test (Zhang & Li, 2021; Kong et al., 2022; Feng et al., 2022; Huang et al., 2023b; Zhang et al., 2023b; a; Feng et al., 2023; Kong et al., 2023) for better structural ability like learning distance between nodes or counting cycles in graph. Some works try to combine the GNN with the transformer. particularly, GraphFormers (Yang et al., 2021) and GROVER (Rong et al., 2020) also insert a GNN layer between consecutive transformer layers for modeling graph inductive bias. Different from us, their transformer layers are randomly initialized and directly tuned on downstream tasks without text.
Appendix E Additional Discussions
E.1 Graph Structure Question Example of LLM
@@
E.2 Theoretical Advantages of GOFA’s Graph Language Encoder
In GOFA, we innovatively integrate GNN layers into LLMs to help LLMs understand graph structures. This approach is theoretically more powerful and suitable for designing GFMs than using pure LLMs. Graph data have unique properties, such as node permutation invariance without fixed ordering (Keriven & Peyré, 2019), making sequential models like LLMs unsuitable for modeling graphs. For a graph with $n$ nodes, the number of possible orders is $n!$, which means sequential models like LLMs will need factorial sample complexity to learn that all of them correspond to the same graph. Existing works therefore use random or heuristic order to represent a graph, resulting in suboptimality and poor generalization on structure-related tasks. One empirical example from the task planning experiments of LLM agents in (Wu et al., 2024) shows that LLMs can only perform well on task graphs with a specific node ordering but cannot maintain this accuracy after nodes are reordered. Methods like LLaGA (Chen et al., 2024) also fall into this category and are suboptimal if the task requires deep structure understanding. Instead, GNNs are a powerful choice widely accepted by the literature for encoding both features and the structure of graphs. They are permutation equivariant to graph order and can learn invariant structure information. Our GOFA , interleaving GNN layers into LLMs, naturally preserves this property. Specifically, for LLM layers, each node is processed individually, which is obvious that we can keep the permutation invariance. As GNN layers are permutation invariant, this conclude that the GOFA is permutation invariant to input graph.
E.3 Advantages of GOFA’s Self-supervised Learning Tasks
Our proposed self-supervised learning tasks are enlightened by existing graph and NLP SSL tasks. However, our tasks are novel compared to existing methods from several perspectives.
In graph SSL, most prior work aims to recover original features or graph structure contrastively or generatively (Liu et al., 2022), using learned embeddings for downstream classifiers. In contrast, our tasks aim to learn embeddings that enable downstream natural language generation. Concretely, our SSL shortest path prediction task requires the model to output multiple actual paths (e.g., node a→node b→node c) between two nodes in text format; this requires a more fine-grained and in-depth model understanding of the graph structure. Regular graph SSL tasks such as link prediction and shortest path distance prediction only cares about a simple objective (binary classification for link and single number for distance). Conversely, our generation-oriented design allows a unified task format to query into the graph from different aspects with different granularity (e.g. shortest path and common neighbor can be incorporated under the same natural language generation framework), whereas traditional SSL tasks inevitability lose much detailed information and may need artificial complicated design to accommodate multiple SSL targets.
SSL in NLP (e.g. next-word-prediction) only takes texts as input. One may think of directly converting graphs to text and doing a similar SSL. However, as many previous works show (Wang et al., 2023), converting graphs directly to texts for LLM generation is ineffective. Hence, we design the sentence completion task directly on the graph to use connection information to help the model attend to correct nodes for subsequent generation.
In summary, our SSL design cares more about training the model to generate any answers in natural language format so that it can accommodate arbitrary tasks, which differs from traditional graph SSL that normally focuses on classification/regression (actually any graph SSL tasks can be incorporated into our natural language generation framework). Compared to NLP SSL, our novel SSL design focuses on sentence completion using neighboring sentence information rather than pure auto-regression, strengthening the model’s power to leverage joint graph-text information.
@@ F.1 General settings
Subgraph sampling: In the GOFA , for node/link/graph-level tasks, the input format is unified as a subgraph task. Namely, for node/link-level tasks, we will select a $k$-hop subgraph surrounding the target nodes as the input graph for the model. We follow a similar subgraph sampling strategy as OFA (Liu et al., 2023a). Specifically, for node-level tasks, we directly sample the $k$-hop subgraph rooted at the target node. Meanwhile, we set a threshold for maximum nodes per hop. If the nodes in a certain hop exceed the threshold, we will randomly sample maximum nodes from all nodes. For link-level tasks, we doing the sampling on both two nodes.
@@
Implementations. Both the GOFA and all baselines are implemented using Python with Pytorch, transformers, and PyG (Fey & Lenssen, 2019) packages.
F.2 Design of pre-training tasks
In this section, we describe the self-supervised pre-training of GOFA . The goal of the pre-training is to let the GOFA model obtain the ability to query graph structure and context but retain the ability to reason about plain text. Specifically, we perform the pre-training task using multiple existing graph datasets, including MAG240M (Hu et al., 2021a), Arxiv (Hu et al., 2021b), Pubmed, Wikikg90mv2 (Hu et al., 2021a), and Ultrachat200k (Ding et al., 2023) datasets. Further, we create another graph dataset called WikiGraph, whose graphs are generated from sentences in the pure textual WikiData dataset (Merity et al., 2022) using LLM. Details about the datasets can be found in Appendix C. We randomly generate a unique node ID (such as [Node A]) for each node in the training sample and append it to the original node text. This ID will serve as a basis for querying nodes in the graph. We design four pre-training tasks: sentence completion, structural understanding, question-answering, and information retrieval tasks. Figure 1 shows an example of each task. We describe the rationale of each task below and leave some implementation details to Appendix F. We also include an additional discussion of the advantages of our designed tasks in Appendix E.3.
Sentence completion task. The objective of the sentence completion task is to train GOFA to reason about the rest of the text in a node given both the existing text and the information in the rest of the graph. Given an input training sample, we randomly select $n$ nodes in the graph as the target nodes. All selected nodes’ texts are split into halves. The first half forms node text $x(v)$, and the second half becomes the target $y$ to generate. The length of the first half will also be randomly determined. Finally, the output representation of these $n$ nodes will be directly input to the decoder (no prompt node will be connected) and we minimize the loss between model decoded text and target $y$. This sentence-completion pre-training task adapts LLMs’ standard ability to the graph context.
Structural understanding tasks. The objective of the structural tasks is to pre-train GOFA to understand basic graph structural properties. In light of this, we design the shortest paths and common neighbors reasoning tasks between nodes. Specifically, For each training subgraph sample, we randomly sample $n$ node pairs as the selected targets. For each selected node pair, we ask the model to compute the shortest path distance between two nodes and output all possible shortest paths between them using the assigned node IDs. Meanwhile, we also ask the model to output the number of common neighbors the two nodes have and the node IDs of their common neighbors. For the structural understanding task, a prompt node $v_{p}$ will connect to both two nodes since our structural tasks need the model to reason about two nodes simultaneously. The text in the prompt node will be the corresponding question. Through these two tasks, the model is expected to gain the ability to identify basic graph structures, which are critical to downstream tasks.
Question answering task. Unlike language corpus, which naturally contains many question-and-answer (QA) pairs, graph data usually only contain objective descriptions of entities. Nevertheless, for the model to be fluid in tasks, we need the model to understand user prompts and be sensitive to different tasks. Hence, we synthesize a QA-chain dataset from Ultrachat200k, as shown in Figure 1. A language QA sequence is converted to a chain graph where nodes with question texts alternate with nodes with answer texts, which are connected by directed edges to represent the conversation order. The last question becomes the text on prompt node $v_{p}$, which is connected to every node in the chain, and the last answer is the target text $y$ (see Figure 1 QA-Chain Task for an example). This QA task provides QA pseudo-graphs missing from the common graph corpus, and we found it critical for enabling the model to be responsive to arbitrary tasks expressed in free-form text questions.
Information retrieval task. For most of the downstream tasks, GOFA requires a prompt node to link to all target nodes in the graph to solve related problems. To enable the prompt node to effectively maintain related information for solving the task in the decoding stage, we design an information retrieval task to realize these goals. Specifically, for each input graph, we randomly select $n$ nodes and we connect a prompt node to these $n$ nodes. Next, the information retrieval task is further divided into two parts: key-to-content and content-to-key. For key-to-content, we provide a node ID (randomly chosen from the selected $n$ nodes) in the prompt node and ask the model to retrieve the text of that node. For the content-to-key task, we provide the content of one node (selected the same as above) in the prompt node and ask the model to return the correct node ID of that node. This task enhances the ability of GOFA to utilize our provided node IDs to retrieve and maintain correct information in the prompt node, which proves useful for many downstream tasks requiring information retrieval.
F.3 Pre-train implementation details of GOFA
@@
Information Retrieval Please output the content of [NODE.A]. Content on [NODE.A].
Given this node content: {node content}, please output the node id. [NODEID].
Dataset and task construction. As we discussed, we designed four different pre-training tasks for GOFA . Here we describe some implementation details about each task and then discuss how we construct each task on each dataset.
For the sentence completion task, the node text is split by the following rule: for each node, if the node text is less than 256 words, we set the maximum left-halve length to be the half of node sentence length. Otherwise, we set it to 128. Next, we randomly choose a length from 0 to maximum left length as the final cut point to cut the sentence into two pieces. For the shortest path task, we ask the model to output both the shortest path distance and all possible shortest paths. Since there may be multiple paths, to ensure the uniqueness of the answer, we first order all paths based on the node ID (the ascending order of alphabets) for nodes in each path and ask the model to learn this order. The construction of common neighbor task is similar. Finally, for information retrieval, given an input graph sample, we randomly select 2 to the number of nodes in the graph to be the target nodes.
For pre-training datasets, we use multiple datasets including MAG240M, Arxiv, Pubmed, Wikikg90mv2, Ultrachat200k, and WikiGraph. For MAG240M, Arxiv, and Pubmed datasets, each training sample is a subgraph sampled around a node. Next, sentence completion, shortest path, and common neighbor tasks are constructed. For each sample, there are 4 complete sentences, 3 shortest path, and 3 common neighbor tasks. We will also construct information retrieval tasks on these datasets. However, to ensure a moderated graph size, the information retrieval task will be constructed separately from the above tasks and also for both key-to-content and content-to-key tasks. For each information retrieval task sample, there will be only one task. For Wikikg90m, each training sample is a subgraph sampled around an edge. In Wikikg90m, we additionally include a link prediction task. That is, for each input graph, we randomly mask $e$ edges and ask the model to recover the content in the edge. For each sample, there are 4 complete sentences, 2 shortest paths, and 2 common neighbor tasks, and 2 link prediction tasks. At the same time, the information retrieval task will also be generated separately. For WikiGraph, each sample is itself a graph. Similar to Wikikg90mv2, each sample consists of 4 complete sentences, 2 shortest paths, 2 common neighbor tasks, and 2 link prediction tasks and information retrieval task will also be generated separately. Finally, for Ultracha200k, we only include question answer task and each sample only contains one task. The detailed task prompts and answer examples are shown in Table 11.
@@
Training details. The initial weight of the LLM compressor and decoder is obtained from ICAE (Ge et al., 2023). The initial weight of all GNN layers is randomly initialized. The value of all gates in the residual connection is set to 0 to ensure the initialized model performs the same as the original language model. During the training, we only tune the GNN layers. For each training epoch, the training corpus includes 500,000 MAG240M samples, 50,000 Arxiv samples, 5,000 PubMed samples, 100,000 Ultrachat200k samples, 80,000 WikiGraph samples, 100,000 Wikikg90mv2 samples. At the meantime, for MAG240M, Arxiv, Pubmed, WikiGraph, and Wikikg90mv2, we will include 10,000 key-to-content and 10,000 content-to-key information retrieval tasks. This resulted in 935,000 samples for each training epoch and we trained the model for 3 epochs. The training is conducted on 8 NVIDIAA100_SXM4_80GB GPUs with DeepSpeed stage 2 (Rajbhandari et al., 2020) parallelism. The detailed training parameters are set the same for both two models and are listed in Table 12. We use AdamW optimizer with $\beta=(0.9,0.95)$. We use a cosine annealing learning rate scheduler, and the minimum learning rate is 10% of the initial learning rate. We restarted the learning rate 2 times on one-third and two-thirds of the training.
@@
Table 13: Prompt examples of GOFA for each training dataset in Zero-shot learning.
@@
MAG240M This is a citation network from microsoft academic graph platform. Nodes represent academic papers and edges represent citation relationship. You are an expert in computer science. You need to choose the correct paper category based on the paper content and its citation network. For example, if the paper [NODEID] {<label_description>, choose <label>;}. What is the most likely paper category for the target paper? Choose from the following: {<label>}.
@@
Pubmed-link This is a co-citation network from the Pubmed platform focusing on diabetes mellitus. Nodes represent academic papers and edges represent two papers that are co-cited by other papers. You are a diabetes mellitus expert tasked with determining whether two given papers [NODEID1] and [NODEID2] are co-cited by another paper based on their content and network characteristics. Evaluate the following criteria: assess whether the topics of the two papers are similar, check if the shortest path distance between the two papers is small, and verify whether the papers have a large number of common neighbors in the citation network. If the answer to most of these questions is Yes, choose Yes; if the answer to most of these questions is No, choose No.
@@
PubMed-node This is a co-citation network from the Pubmed platform focusing on diabetes mellitus. Nodes represent academic papers and edges represent two papers that are co-cited by other papers. You are an expert on diabetes mellitus. You need to choose the correct paper category based on the paper content and its co-citation network. For example, if the paper [NODEID] {<label_description>, choose <label>;}. What is the most likely paper category for the target paper? Choose from the following: {<label>}.
@@
Wikikg90m This is a graph extracted from the entire Wikidata knowledge base. You are an expert in knowledge graph reasoning. You need to choose the correct relation type between two target entities based on their existing relations. For example, if two relations involve {<label_description>, choose <label>;}. What is the relationship between two target entities? Choose from the following list: {<label>}."
Arxiv_node This is a citation network from Arxiv platform focusing on the computer science area. Nodes represent academic papers and edges represent citation relationships. You are an expert in computer science. You need to choose the correct paper category based on the paper content and its citation network. For example, if the paper [NODEID] {<label_description>, choose <label>;}. What is the most likely paper category for the target paper? Choose from the following: {<label>}.
Arxiv_link This is a citation network from Arxiv platform focusing on the computer science area. Nodes represent academic papers and edges represent citation relationships. You are a computer science expert tasked with determining whether two given papers [NODEID1] and [NODEID2] are co-cited by another paper based on their content and network characteristics. Evaluate the following criteria: assess whether the topics of the two papers are similar, check if the shortest path distance between the two papers is small, and verify whether the papers have a large number of common neighbors in the citation network. If the answer to most of these questions is Yes, choose Yes; if the answer to most of these questions is No, choose No.
Table 14: Prompt examples of GOFA for each evaluation dataset in Zero-shot learning.
Dataset Prompt
Cora-node This is a co-citation network focusing on artificial intelligence, nodes represent academic papers and edges represent two papers that are co-cited by other papers. You are an expert in computer science. You need to choose the correct paper category based on the paper content and its co-citation network. For example, if the paper [NODEID] {<label_description>, choose <label>;}. What is the most likely paper category for the target paper? Choose from the following: {<label>}.
Cora-link This is a co-citation network focusing on artificial intelligence, nodes represent academic papers, and edges represent two papers that are co-cited by other papers. You are a computer science expert tasked with determining whether two given papers are co-cited by another paper based on their content and network characteristics. Evaluate the following criteria: assess whether the topics of the two papers are similar, check if the shortest path distance between the two papers is small, and verify whether the papers have a large number of common neighbors in the citation network. If the answer to most of these questions is Yes, choose Yes; if the answer to most of these questions is No, choose No.
WikiCS This is a Wikipedia graph focusing on computer science. Nodes represent Wikipedia terms and edges represent two terms that have hyperlinks. You are an expert in computer science. You need to choose the correct category of Wikipedia term based on the term content. For example, if the term [NODEID] {<label_description>, choose <label>;}. What is the most like category for this Wikipedia entry? Choose from the following: {<label>}.
Products This is a co-purchase network from the Amazon platform. Nodes represent the products sold on Amazon and edges represent two products that are co-purchased together. For example, if the product [NODEID] {<label_description>, choose <label>;}. What is the most like category for this product? Choose from the following: {<label>}.
FB15K237 This is a knowledge graph from the FreeBase. Nodes represent knowledge entities and edges represent relations between two entities. You are an expert in knowledge graph reasoning. You need to choose the correct relation type between two target entities based on their existing relations. For example, if two relations {<label_description>, choose <label>;}. What is the relationship between two target entities? Choose from the following list: {<label>}."
@@
F.4 Zero-shot learning
Setting. For the zero-shot learning, we select Cora-link, Cora-node, WikiCS, Products, ExplaGraphs, and SceneGraphs as evaluation datasets. For all datasets, we directly evaluate baselines and GOFA on the test set.
Baseline Details: We compare the performance of GOFA with two categories of baseline methods. The first category includes models that directly utilize large language models (LLMs). For this, we select Llama2-7B and Mistral-7B (Jiang et al., 2023) as baselines. We input the content of all target nodes into these pre-trained models and concatenate the same prompt used in GOFA for evaluation. The second category consists of Graph LLM models that have zero-shot ability. We include OFA (Liu et al., 2023a), GraphGPT (Tang et al., 2023), UniGraph (He & Hooi, 2024), ZeroG (Li et al., 2024), and LLaGA (Chen et al., 2024) as baselines. For OFA, we extend the datasets by adding Products and ExplaGraphs and follow the original source code to train the model on Arxiv and FB15K237 for 30 epochs, using Llama2-7B as the embedding model. All other settings remain consistent with the default OFA configuration, and we report the test performance accordingly. For GraphGPT, we use the results reported in the LLaGA paper. For UniGraph, we use the results from the original paper. For ZeroG, we use the results in UniGraph paper. For LLaGA, we rerun the source code, adapting the settings of ways to align with our experimental setup.